A 36 ohm resistor and an 18 ohm resistor are connected in series across a 15V battery. What is the voltage across a) the 36 ohm resistor. b) the 18 ohm resistor?


1
Expert's answer
2021-06-23T13:09:13-0400

Gives

"R_1=36\Omega,R_2=18\Omega"

V=15V

Both resistance are series combination

"R_{net}=R_1+R_2"


"R_{net}=36+18=54\Omega"


"I_{net}=\frac{V}{R}=\frac{

15}{54}"



"I_{net}=0.277A"

"V_1=IR_1=0.277\times36=9.972V"



"V_2=IR_2=0.277\times18=4.986V"